OpenStudy (anonymous):

Which is the solution of the system x + y = -2 2x - 3y = -9 A. (-4,2) B. (-3,1) C. (-3,-5) D. (4,-6)

terenzreignz (terenzreignz):

What method are you used to using, in solving systems?

OpenStudy (anonymous):

I dunno

terenzreignz (terenzreignz):

Let's stick with the basic substitution method then. Let's look at the equation x + y = -2 Could you solve for y here? By solve for y, I mean, express y in terms of x. Or, in other words, bring all terms without y to the right side...

OpenStudy (anonymous):

I'm still confused.

terenzreignz (terenzreignz):

Well then, x + y = -2 What happens when you subtract x from both sides?

OpenStudy (anonymous):

I don't know. I'm retarded with this stuff.

OpenStudy (cwrw238):

x + y = -2 subtract x from both sides: x - x + y = -2 - x y = -2 - x the x has disappeared on left hand side and we have y in terms of x do you follow that?

OpenStudy (cwrw238):

the idea is to get y on its own on one side of the '=' and we achieved this by getting rid of the x on lefthand side. an equation is like a chemical balance whatever you do to one side you have to do to the other to maintain the balance we subtracted x from LHS so we had to subtract from RHS

OpenStudy (anonymous):

hey the easy method which saves time when we donno the process is direct substitution this may help some times the correct ans is option B.(-3,1)